Write the equation in standard form for the parabola with the vertex (0,0) and directrix y=-8
Answer by addingup(3677)

The standard form is: (x-h)^2 = 4p(y-k)
You say that the distance from the vertex to the directrix is p = -8
Substitute for h,k and p to get:
(x-0)^2 = -20(y-0)
-----
x^2 = -32y
y = (-1/32)x^2